OpenStudy (anonymous):

1.Suppose that in one week, your company takes in $2,220 in sales at one of its cell phone stores The price breakdown per phone is as follows: Cell phone model A4: $35 Smartphone Z20: $50 Suppose that a total of 51 phones were sold.

OpenStudy (anonymous):

cell phone model A4 is 22 phones and smartphone Z20 is 29

OpenStudy (anonymous):

Take what you know, a + z = 51, to solve for z it becomes z = 51-a Then to figure out the next part, 35a + 50z = 2220 Replace z with 51-a... it becomes 35a + 50 (51-a) = 2220